OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1=-7 and a18=95

OpenStudy (anonymous):

@SolomonZelman

hartnn (hartnn):

Do you know the formula for n'th term of arithmetic sequence ?

OpenStudy (anonymous):

no not really

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ a_{18}=a_{1} +d(18-1) }\)

OpenStudy (solomonzelman):

and you know the a18 and a1 find the difference then find the 35th term

OpenStudy (anonymous):

do you just plug in the other numbers?

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf95=-7+d(18-1) }\) solve for d

OpenStudy (anonymous):

k

OpenStudy (aum):

The nth term of an arithmetic sequence is: \(\Large a_n = a_1 + (n-1)d\) You are given the first term and the eighteenth term. From these two pieces of information you can find the common difference 'd' Plug the numbers into the above formula and find d. Then use the same formula to find the 35th term (by setting n = 35)

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

is it 6?

OpenStudy (aum):

Yes, d = 6.

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

203 197 168 163

OpenStudy (aum):

You have to find the 35th term. The nth term of an arithmetic sequence is: \(\Large a_n = a_1 + (n-1)d\) Plug in the numbers: \(\Large a_n = -7 + (n-1)*6\) Find the 35th term by setting n = 35: \(\Large a_{35} = -7 + (35-1)*6 = ?\)

OpenStudy (anonymous):

197

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

thanks for the help

OpenStudy (aum):

You are welcome.